# FreeCap: Hybrid Calibration-Free Motion Capture in Open Environments

### v1

@@ Introduction
Camera-based methods (Belagiannis and Zisserman 2016; Miezal, Taetz, and Bleser 2017; Rajasegaran et al. 2022; Shin et al. 2024) can effectively capture the accurate human local pose by image texture information but lack depth information, and depth camera-based methods (Sridhar et al. 2015; Wei, Zhang, and Chai 2012; Yu et al. 2017, 2019; Zheng et al. 2018) suffer from lighting conditions and the limited sensor range. LiDAR is widely used in robotics and autonomous driving (Zhu et al. 2021, 2020; Cong et al. 2022; Xu et al. 2023) for perceiving depth information in large-scale scenes. Previous LiDAR-based methods (Li et al. 2022; Fan et al. 2023b; Ren et al. 2024) can estimate the global human pose in large-scale scenes by accurate depth information. However, the sparsity of long-distance point clouds and the lack of texture information can lead to inaccuracies in predictions.
To leverage both image texture and LiDAR depth information for more accurate global human motion predictions in open environments. The previous method(Cong et al. 2023; Fan et al. 2023a) utilizes calibrated multiple cameras and LiDARs for involving the fusion of raw data, which can lead to difficulties in unifying data distribution and reduced generalization ability. In contrast, as shown in Fig. LABEL:fig:teaser, our system consists of the single LiDAR and any number of moving cameras. The LiDAR captures extensive global depth information, while the moving cameras provide detailed texture information in local areas. This more flexible and expandable setup enables accurate multi-person motion estimation within a unified world coordinate system, making it suitable for diverse applications.
@@
Early motion capture system reconstructs human mesh relays on dense markers  (Loper, Mahmood, and Black 2014; Park and Hodgins 2008; Song and Godøy 2016; Raskar et al. 2007; Xu et al. 2019) attach in human body. It enables the capture of high-quality human motions and is wildly applied in industry. However, the system is costly and limited in indoor scenes with good lighting conditions. The freer system uses inertial measurement units (Yi, Zhou, and Xu 2021; Yi et al. 2022; Ren et al. 2023) for capturing orientations and accelerations of human key bones which reflect the human motion representation based on SMPL model. It can work in relatively open scenarios, but it suffers from drift and can be affected by the magnetic field. Moreover, the above methods require actors to wear sensors, causing constrained human motions.
@@
To make the mocap method applied in daily usage, the markerless methods (Liu et al. 2013; Rhodin et al. 2016; Kanazawa et al. 2018) use cameras only has made great progress. The multi-view camera-based methods (Malleson, Collomosse, and Hilton 2019; Zhang et al. 2020; Malleson, Collomosse, and Hilton 2020) predict accurate human motions based on optimization from full perspective motion information. Though the performers do not need to wear markers, the multi-camera studio needs complex deployment in closed areas. The monocular-based method (Mehta et al. 2017; Sapp and Taskar 2013) only requires a single camera, it is light-weight while the performer is free to move, but a single camera can not perceive depth information and the sensor range is narrow. Recently, some methods (Shin et al. 2024; Rajasegaran et al. 2022), begin to pay attention to estimating the global human trajectory from a monocular dynamic camera. However, the trajectory is relative to a beginning frame and accumulated errors occur, which means that methods cannot infer long-term data and cannot predict multi-person global motions in a unified world coordinate system.
@@
In order to estimate the global human motions in large-scale open environments, various LiDAR-based mocap methods have been proposed. LiDARCap (Li et al. 2022) first proposes a single LiDAR-based human motion capture method, but it only estimates the local pose based on the graph-based framework. LiDAR-HMR (Fan et al. 2023b) utilizes the point cloud geometric information to reconstruct the human mesh, but it can not provide the template-based motion representations, the scope of application is limited. PointHPS (Cai et al. 2023) provides a cascaded network to predict the human pose and shape, but the network framework requires dense points in close range. LiveHPS (Ren et al. 2024) estimate the full SMPL parameters, including the human local pose, body shape, and global translation, fully utilizing the depth information of point cloud. However, the sparsity of data and the lack of texture information also lead to inaccurate predictions.
@@
Considering human motions are coherent over time, we employ temporal interactions to leverage the sequential nature of the data to capture temporal dependencies and correlations, thereby improving the overall results. We use multi-head self-attention layer to capture this temporal relationship and get the encoded feature $f_{3d}$ and $f_{2d}$.
Subsequently, we interact the features from the arbitrary two modalities(or two sensor features) by designing a bidirectional cross-attention module and get the $f_{3d}$ and $f_{2d}$ from two distinct modalities.
@@
We build our framework on PyTorch 2.0.0 and CUDA 11.8 and run the whole process on a server equipped with an Intel(R) Xeon(R) 444 E5-2678 CPU and 8 NVIDIA RTX3090 GPUs. For PCM, we set $\delta$ to 100, $\lambda_{0}$ to 0.1 and $n_{iter}$ to 2 to get a stable matching. During the training of SPO, we train the network over 500 epochs with batch size of 32 and sequence length of 32, using an initial learning rate of $10^{-4}$, and AdamW optimizer with weight decay of $10^{-4}$. We set $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=0.01$ throughout our experiment. As for SMPL solver, we set $\lambda_{4}=1$, $\lambda_{5}=0.2$, $\lambda_{6}=10$ and $\lambda_{7}=1$. For the experiment, we consistently employed the SURREAL (Varol et al. 2017) dataset for pre-training throughout the entirety of our experimental procedure, mirroring the approach adopted by LiveHPS. Following this, we pre-trained the WHAM method on the SURREAL dataset and subsequently refined it through fine-tuning on both the Freemotion (Ren et al. 2024) and Human-M3(Fan et al. 2023a) datasets.
@@
Matching Strategy. To further demonstrate the accuracy and efficiency of our matching method, we conduct a detailed ablation study on the PCM, as shown in Tab. 3. The “KPs" represents we calculate all 2D key points and all 3D key points in every possible matching pair, which requires the time complex is $\mathcal{O}(n^{n})$. The “KP" represents we random select 3D key points of one person and calculate with all 2D key points, the time complex is $\mathcal{O}(n)$ but the accuracy decrease. “Pose" represents we only use the estimated body pose for single frame matching and “Pose(T)" represents we use the sequential body pose for sequence matching. “Pose & KPs" represents our matching method for single frame matching without temporal information. Our method achieves the highest accuracy while balancing efficiency. Fig. 3 demonstrates our method’s performance in extremely complex scenes with ten individuals, our method works well even if some persons in the camera are missing.
@@
KPs KP Pose Pose & KPs Pose(T) Pose(T) & KPs
Accuracy(%$\uparrow$ 89.88 77.07 81.59 85.87 96.01 98.10
[…]
Expandable Sensor. As shown in Tab. 4, we evaluate in FreeMotion with different settings. Since the narrow sensor range of camera, in some extreme cases, the camera cannot capture complete human information. We reorganize the testing dataset of FreeMotion by using cameras with poor viewing perspective. As the number of cameras increases the human motion information obtained becomes more complete, and we can achieve better performance, which proves that our sensor-expandable pose optimizer can fully utilize the effective information of all sensors and can work even when losing one modality data.
Table 4: Ablation study for our expandable sensors in FreeMotion.
[…]
1 0 59.45/73.04 100.67/108.82 12.87 6.09 4.93
0 1 81.00/96.49 119.77/129.83 15.41 7.38 6.79
1 1 57.66/70.52 99.64/107.31 11.82 6.06 5.21
1 2 56.07/68.03 98.32/105.54 11.66 5.82 3.92
1 3 54.21/66.51 96.50/103.52 11.40 5.92 4.22

### v2

@@ Abstract
Figure 1: Visualization of our FreeCap in a real-time captured scenario. Our settings include a single LiDAR and four cameras. Camera-1 follows the running person, camera-2 surrounds two people playing soccer, camera-3 focuses on the main person playing frisbee and camera-4 captures three persons. We zoom in some cases to the right.
@@ Introduction
Camera-based methods (Belagiannis and Zisserman 2016; Miezal, Taetz, and Bleser 2017; Rajasegaran et al. 2022; Shin et al. 2024) can effectively capture the accurate human local pose by image texture information but lack depth information, and depth camera-based methods (Sridhar et al. 2015; Wei, Zhang, and Chai 2012; Yu et al. 2017, 2019; Zheng et al. 2018) suffer from lighting conditions and the limited sensor range. LiDAR is widely used in robotics and autonomous driving (Zhu et al. 2021, 2020; Cong et al. 2022; Xu et al. 2023; Lu et al. 2023) for perceiving depth information in large-scale scenes. Previous LiDAR-based methods (Li et al. 2022; Fan et al. 2023b; Ren et al. 2024) can estimate the global human pose in large-scale scenes by accurate depth information. However, the sparsity of long-distance point clouds and the lack of texture information can lead to inaccuracies in predictions.
To leverage both image texture and LiDAR depth information for more accurate global human motion predictions in open environments. The previous method(Cong et al. 2023; Fan et al. 2023a) utilizes calibrated multiple cameras and LiDARs for involving the fusion of raw data, which can lead to difficulties in unifying data distribution and reduced generalization ability. In contrast, as shown in Fig. 1, our system consists of the single LiDAR and any number of moving cameras. The LiDAR captures extensive global depth information, while the moving cameras provide detailed texture information in local areas. This more flexible and expandable setup enables accurate multi-person motion estimation within a unified world coordinate system, making it suitable for diverse applications.
@@
Early motion capture system reconstructs human mesh relays on dense markers  (Loper, Mahmood, and Black 2014; Park and Hodgins 2008; Song and Godøy 2016; Raskar et al. 2007; Xu et al. 2019) attach in human body. It enables the capture of high-quality human motions and is wildly applied in industry. However, system is costly and limited in indoor scenes with good lighting conditions. The freer system uses inertial measurement units (Yi, Zhou, and Xu 2021; Yi et al. 2022; Ren et al. 2023) for capturing orientations and accelerations of human key bones which reflect the human motion representation based on SMPL model. It can work in relatively open scenarios, but it suffers from drift and can be affected by magnetic field. Moreover, above methods require actors to wear sensors, causing constrained human motions.
@@
To make mocap method applied in daily usage, the markerless methods (Liu et al. 2013; Rhodin et al. 2016; Kanazawa et al. 2018) use cameras only has made great progress. The multi-view camera-based methods (Malleson, Collomosse, and Hilton 2019; Zhang et al. 2020; Malleson, Collomosse, and Hilton 2020) predict accurate human motions based on optimization from full perspective motion information. Though performers do not need to wear markers, the multi-camera studio needs complex deployment in closed areas. The monocular-based method (Mehta et al. 2017; Sapp and Taskar 2013) only requires a single camera, it is light-weight while the performer is free to move, but a single camera can not perceive depth information and the sensor range is narrow. Recently, some methods (Shin et al. 2024; Rajasegaran et al. 2022), begin to pay attention to estimating the global human trajectory from a monocular dynamic camera. However, the trajectory is relative to a beginning frame and accumulated errors occur, which means that methods cannot infer long-term data and cannot predict multi-person global motions in a unified world coordinate system.
@@
In order to estimate the global human motions in large-scale open environments, various LiDAR-based mocap methods have been proposed. LiDARCap (Li et al. 2022) first proposes a single LiDAR-based human motion capture method, but it only estimates the local pose based on the graph-based framework. LiDAR-HMR (Fan et al. 2023b) utilizes the point cloud geometric information to reconstruct the human mesh, but it can not provide the template-based motion representations, the scope of application is limited. PointHPS (Cai et al. 2023) provides a cascaded network to predict the human pose and shape, but the network framework requires dense points in close range. LiveHPS and LiveHPS++ (Ren et al. 2024, 2025) estimate the full SMPL parameters, including the human local pose, body shape, and global translation, fully utilizing the depth information of point cloud. However, the sparsity of data and the lack of texture information also lead to inaccurate predictions.
@@
Considering human motions are coherent over time, we employ temporal interactions to leverage the sequential nature of the data to capture temporal dependencies and correlations, thereby improving the overall results. We use multi-head self-attention layer to capture this temporal relationship and get the encoded feature $f_{3d}$ and $f_{2d}$. Subsequently, we interact the features from the arbitrary two modalities(or two sensor features) by designing a bidirectional cross-attention module and get the $f_{3d}$ and $f_{2d}$ from two distinct modalities.
@@
We build our framework on PyTorch 2.0.0 and CUDA 11.8 and run the whole process on a server equipped with an Intel(R) Xeon(R) 444 E5-2678 CPU and 8 NVIDIA RTX3090 GPUs. For PCM, we set $\delta$ to 100, $\lambda_{0}$ to 0.1 and $n_{iter}$ to 2 to get a stable matching. During the training of SPO, we train the network over 500 epochs with batch size of 32 and sequence length of 32, using an initial learning rate of $10^{-4}$, and AdamW optimizer with weight decay of $10^{-4}$. We set $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=0.01$ throughout our experiment. As for SMPL solver, we set $\lambda_{4}=1$, $\lambda_{5}=0.2$, $\lambda_{6}=10$ and $\lambda_{7}=1$. For the experiment, we consistently employed the SURREAL (Varol et al. 2017) dataset for pre-training throughout our experimental procedure, mirroring the approach adopted by LiveHPS. Following this, we pre-train the WHAM method on the SURREAL dataset and subsequently refine it through fine-tuning on both the Freemotion (Ren et al. 2024) and Human-M3(Fan et al. 2023a) datasets.
@@
Matching Strategy. To further demonstrate the accuracy and efficiency of our matching method, we conduct a detailed ablation study on the PCM, as shown in Tab. 3. The “KPs" represents we calculate all 2D key points and all 3D key points in every possible matching pair, which requires the time complex is $\mathcal{O}(n^{n})$. The “KP" represents we random select 3D key points of one person and calculate with all 2D key points, the time complex is $\mathcal{O}(n)$ but the accuracy decrease. “Pose" represents we only use the estimated body pose for single frame matching and “P & T" represents we use the sequential body pose for sequence matching. “P & K" represents our matching method for single frame matching without temporal information. Our method achieves the highest accuracy while balancing efficiency. Fig. 4 demonstrates our method’s performance in extremely complex scenes with ten individuals, our method works well even if some persons in the camera are missing.
KPs KP Pose P & K P & T P & T & K
Acc.(%$\uparrow$ 89.88 77.07 81.59 85.87 96.01 98.10
[…]
@@ Conclusion
Acknowledgments
This work was supported by NSFC (No.62206173), Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (KLIP-HuMaCo).